# Shared Hardships Strengthen Bonds: Negative Shocks, Embeddedness and Employee Retention

### v1

@@ 3 Theory and Hypotheses
Driver and employer interests are not aligned for all shocks. Pay and home time provide important counter-examples (Burks et al., 2023). Piece-rate pay per mile implies that any pay increase for drivers reduces the residual for the firm; moreover, pay increases have not been shown to increase productivity in this industry (see, for example,(Belzer & Sedo, 2018) on income targeting). More home-time for drivers, means less capacity for the firm and lowered capital utilization when drivers take vehicles home (Burks et al., 2023), Conversations between employers and employees about pay and home time are thus more likely to be acrimonious and less likely to result in fruitful resolution. The resultant interest divergence reduces embeddedness, increasing turnover. Formally, we have H2:
Hypothesis 2.
Shocks which expose interest-divergence increase turnover.
@@ 4.1 Data
Our empirical analysis uses communication records from a human resource outsourcing company specializing in retention programs for trucking companies. The core of their programs is making weekly (or longer interval) phone calls to truck drivers to identify and record their “issues”– “shocks” in the terminology of the unfolding model of turnover– and then pass them to their customers (i.e., trucking companies) to help them to be proactive in driver retention. Specifically, the retention programs typically last for three to six months for each driver and during the program, the agents of the company record a brief summary of their conversation after each call and tag it with proprietary classification codes for the issues that they identify during the conversation (hereafter issue codes). Some examples of the communication notes are below:
@@
The classification system includes more than two hundreds issue codes; we re-classify these codes into fourteen major categories: Home-time, Pay, Equipment, Safety/Ethics, Training, Team, Operations, Disrespect, Company, Covid, Waiting, Personal, Health, Others. [^3] Due to an non-disclosure agreement with the human resource company, the full list of issue codes and their definitions can’t be shared here. Instead, we include some examples for each issue category in Appendix A. In our main analysis, we further aggregate the categories into two: “Aligned,” where employee-employer interests are aligned, including Equipment, Operations and Waiting, and “Misaligned,” where their interests are opposed, which includes all other categories.
@@
We merge these communication records with driver employment data to create survival-time data suitable for duration analysis. The employment data includes hire and termination dates, from which we calculate employment duration (in weeks) and termination status at the time of each communication record. For drivers whose employment was not terminated during the program, no termination date is observed and their employment spells are treated as censored (i.e., right-censoring). To create a more homogeneous sample, we exclude non-voluntary termination cases (i.e., disqualified quit) and employment spells longer than 100 weeks. [^4] Our results in the next sections are robust to this sample selection. The data is then merged with employer data such as carrier-level scores related to unsafe driving, hours-of-service compliance, vehicle maintenance, misuse of controlled Substances/alcohol from Safety Measurement System (SMS) of the Federal Motor Carrier Safety Administration.
@@
The duration and termination status variables, along with the communication records, are the key variables in our duration analysis. Our data is multi-record-per-subject survival data, where individual drivers are observed multiple times with different communication records over their employment spell. This provides our analysis additional strength because it allows for time-varying regressors that can account for the entire history of covariates. This feature is particularly useful for analyzing the effects of lagged variables.
@@
Our final sample includes 466,236 communication records (= total observations) of 45,873 employment spells from 21 carriers over 2019 - 2022. Most spells are from large carriers, with the largest carrier accounting for 60% and other seven carriers accounting for 32.6% of the sample. [^5] Employment concentration at larger firms is typical in this industry. While the majority of firms are small, the power-law distribution of firms size dictates that the majority of workers work for larger firms (Balthrop, 2021) Later years includes more observations: 16,882 (2019), 71,839 (2020), 158,221 (2021), 219,294 (2022). [^6] These are the numbers of communication records from the spells initiated in a given year. The sample includes four different retention programs, four call-intervals, and fifteen agents, where two programs account for 97.8% of the sample and most observations are based on weekly or bi-weekly phone calls (99.8%). As we shall see later, our duration analysis controls for heterogeneity across carriers, years, months, programs, call-intervals and agents using fixed effects and flexible hazard functions. [^7] In practice, the retention programs and call-intervals are determined by carriers, and in the data, program type and call-interval do not vary much within carriers (i.e., each carrier applies the same program and call-interval to all their drivers). This indicates that controlling for carrier-level heterogeneity is particularly important and an effective way to control for overall heterogeneity in our data.
@@
Next, we provide Kaplan-Meier (KM) survival curves in Figure 1 to summarize some important features in the duration data. The KM survival curves non-parametrically estimate the survival rates at each employment week (i.e., the proportion of employment spells that were not terminated after a given number of weeks). The shaded areas represent 95% confidence intervals. We estimate the curves by carrier and year to highlight some of important heterogeneity in the data. The carriers are ranked in terms of the number of employment spells and their rankings are denoted by “Top-.”
@@
The survival curves demonstrate the high turnover rate in trucking industry. Top1 and Top2-8 carriers’ survival rates are below 0.5 after 24 weeks, indicating that more than 50% of the drivers in these carriers quit their job within six months. A significant heterogeneity across carriers is observed. The survival rate of Top2 carrier is significantly higher than the other carriers, with the gap widening as the employment duration increases. The heterogeneity across years is also interesting. The survival curve of year 2019 flattens around fifteenth week, implying that driver turnover significantly decreases after the period. However, after 2020, turnover worses, even for long-duration employees.
@@ 4.2 Empirical Model: Cox Proportional Hazard Model
We use fixed effects and flexible hazard functions to control for heterogeneity across carriers, years, months, programs, call-intervals and agents. In all specifications, we include fixed effects for year, month, retention program, call-interval and agent. Since much of the retention program’s details (length, call-interval, etc) are determined by carriers, and carrier characteristics are arguably the most important factors in driver’s quit decision, we employ three different approaches to control for heterogeneity across carriers: fixed effects, carrier-level variables from the SMS, and carrier-specific baseline hazard functions. In the result tables in the next sections, Model 1 and 2 includes carrier fixed effects and carrier-level variables, respectively, and Model 3 allows for carrier-specific baseline hazard functions. [^12] The partial likelihood estimation of the Cox PH model calculates the likelihood of each (observed) termination event by the hazard rate of the spell terminated at the event divided by the sum of the hazard rates of the terminated and non-terminated (i.e., censored) spells existed at the event. The division in the calculation cancels out the baseline hazard function due to the specification in (1), which allows us to estimate $\beta$ without specifying the baseline hazard function. When group-specific baseline hazard functions are allowed, the ”partial” likelihoods are calculated by group.
@@ 5.1 Effects of Issue Types on Driver Turnover
To examine our hypotheses, Table 2 reports the coefficient estimates for “Aligned” and “Misaligned” issues in the Cox PH model. [^13] Remember that the estimate for each issue type is the average of the estimates for its original issue dummy and four lags, so it includes the dynamic effects. As defined in Section 4.1, “Aligned,” where employee-employer interests are aligned, includes Equipment, Operations and Waiting, and “Misaligned,” where their interests are opposed, includes all other categories. Coefficients estimates are stable across specifications.
@@
A stark contrast between the two types of issues is observed: all other things being equal, “Aligned” issues on average decrease driver turnover, whereas “Misaligned” issues increase it. “h.r.” are hazard ratios, which helps the reader quickly appreciate the magnitude of the impact. For example, in Model 1, the hazard ratio of Aligned is $-0.29$, indicates that experiencing a shock that highlights aligned interests decreases the hazard (i.e., instantaneous probability of leaving the job) by 25% over the baseline hazard. On the contrary, the hazard ratio for Misaligned issues is $0.32$, indicates having an issue like Home-time, Pay, etc increases the hazard by nearly 38%. The somewhat counter-intuitive result here is that Aligned issues lower driver turnover, even though they are still negative events that drivers experienced. For example, a driver that experiences problems with having to wait at a loading dock is more likely to remain at the motor carrier than a driver that experiences no wait-time shock.
We argue that this is because the Aligned issues provided opportunities for both drivers and employers to communicate with positive messages and confirm their mutual interests, thus increasing the employee’s embeddedness. When the driver experiences operational difficulty, equipment failure, or extended wait times, these shocks also harm the firm by directly reducing worker and equipment utilization. A worker detained at a loading dock, or broken down on the roadside, is not transporting freight. In these cases the motor carrier as an organization has incentive to sympathize with workers, which may result in strengthened relational linkages between frontline workers (drivers) and their support teams. Beyond words, the firm is also likely to take action to remedy the problem. Whether through words (which we observe) or actions (unobserved) these Aligned shocks highlight the firm and workers’ common purpose, while potentially revealing the firm’s support infrastructure available to the worker.
We further disaggregate these results in Table 3, which reports the coefficient estimates for fifteen major issue types in the Cox PH model. [^14] More detailed results including the estimates of individual lag terms are available upon request.
@@
As with our main results, coefficient estimates in Table 3 are not sensitive to specification, whether carrier fixed effects (Model 1), carrier-level variables (Model 2) or carrier-specific baseline hazards (Model 3) are included. Categories most strongly associated with increased turnover include Covid and Personal issues. While Covid is a strong indicator of potential exit, it is a comparatively rare issue (only 224 instances), even though our sample period covers the onset of the pandemic. Personal issues are more common, and have stronger effects on turnover. This category encompasses issues related to family, which supports previous research establishing loneliness and family disruption as one of the most important reasons for truck driver turnover (Williams et al., 2017). Even the smaller (but positive) coefficients have managerially meaningful effects on hazard: ethics complaints increase the baseline hazard rate by 25% (model 1). In contrast, the aligned issues, including Equipment, Operations, and Waiting, all result in strongly reduced hazard rates.
@@
A potential concern, however, is that there may be unobserved driver types with different hazard rates and these types are correlated with the type of issues they reported. For example, there may be “attached” and “detached” drivers with low and high turnover rates, and the former report aligned issues more frequently than the latter, simply because they are more attached to their company. Then, the signs of the coefficients on Aligned and Misaligned issues are simply a reflection of the hazard rate difference between the two types, not the effects of the two different types of issues. Said differently, it is possible to imagine a driver experiencing an equipment issue but not reporting it precisely because the driver expects the issue will go unresolved. Such a disillusioned driver can be expected to have a higher turnover rate; importantly, this type of detached driver will fall in the base-case “no issue” category, biasing base-case turnover rates upward.
To address this potential endogeneity issue, we employ two approaches. First, we create a variable quantifying the responsiveness of each driver to the human resource company’s calls. Responsiveness is an important revealed characteristic that can be correlated with driver types. We calculate the proportion of the calls for each driver tagged with “voice mail,” “no answer,” etc and include it into the duration models. The call answer rate is quite uniformly distributed across drivers, with quartiles 0.20, 0.42, and 0.64. The results are reported in Table B1 in Appendix B, where the estimates largely remain the same.
@@
Next, we focus on the sub-population of communicators. That is, rather than making comparison to a reference group that may under-report the shock, we make comparisons between groups that report multiple shocks. First, we dis-aggregate the Aligned category into “Broken-equipment” and “Other aligned,” where “Broken-equipment” includes issues related to “broken/missing components” only, which is arguably less subject to the under-reporting problem. Second, we re-classify the issues into new categories that account for joint-distribution of Aligned and Misaligned issues: Broken-equipment only (B-only), Other aligned only (OA-only), Misaligned only (MA-only) and their combinations. [^15] Note that a communication record can be tagged with multiple codes in difference categories. Communication records counted multiple times for the fourteen categories in Table 1 can be counted fewer times for the new categories in Table 4, resulting in different total numbers in the tables.
@@
We note that, if other covariates are ignored, the coefficient on “Aligned” in Table 2 is calculated by a combination of two hazard rate differences: a hazard rate difference between the case experiencing Aligned issues only and the case experiencing no issues, and another hazard rate difference between the case experiencing both Aligned and Misaligned issues and the case experiencing Misaligned issues only. Roughly speaking, the first difference compares drivers reporting Aligned issues only and drivers who were silent, whereas the second difference compares drivers who report at least one of the two types of issues. Because the second comparison is among “communicators,” it is arguably less subject to the under-reporting issue. The new categories allow us to identify these dis-aggregated effects.
@@
Table 4 include the distribution of the issues under the new categories, and Table 5 reports the coefficient estimates for the new categories. The coefficient on “Broken & Misaligned” subtracted by the coefficient on “Misaligned” calculates the effect of Aligned issues that is less subject to the under-reporting issue as argued above, which is -0.14 in Model 1 (i.e., specification including carrier fixed effects), with standard error 0.08, t-statistic -1.78 and p-value 0.07.161616The standard error is calculated using the delta method. The results from Model 2 and 3 specifications are qualitatively the same. The p-value is above 5%, but the result overall indicates that relative to the population experiencing only Mis-aligned shocks, a driver experiencing both an Aligned and Mis-aligned shock has a lower hazard rate. The striking result that two negative shocks are better than one shock provides strong support for our hypotheses, at least in this sub-population.

### v2

@@ 3.2 Hypothesis Development
Driver and employer interests are not aligned for most of the listed shocks. Pay and home time provide important counter-examples (Burks et al., 2023). Piece-rate pay per mile implies that any pay increase for drivers reduces the residual for the firm. Pay increases have not been shown to increase driver productivity in this industry (see, for example, Belzer & Sedo, 2018, on income targeting) so there are no output increases–either. Thus, pay-rates are a pure cost which the firm has incentive to minimize. Similarly, more home-time for drivers, means less capacity for the firm and lowered capital utilization when drivers take vehicles home (Burks et al., 2023). Conversations between employers and employees about pay and home time are thus more likely to be acrimonious and less likely to result in fruitful resolution. The resultant interest divergence reduces embeddedness, increasing turnover.
Driver and Employer interests are misaligned for the majority of shocks we study, including home-time, pay, training, teams, disrespect, company policy, Covid, personal matters, and health. In none of these cases do employee issues attach to the fundamental business operations determining firm profitability, except through the effect on employee retention.
The safety/ethics category of shocks present an interesting edge case in terms of interest alignment. For severe safety issues potentially resulting in a crash, interests are certainly aligned, not least because of the potential legal liabilities for the firm, and injuries to drivers. However, our data sample is made up of comparatively safe firms, so we do not observe many of these types of issues. Instead, reported ethical transgressions, such as dishonesty in recruiting, are comparatively more frequent and clearly misaligned with respect to employee and employer interests. We therefore provisionally classify safety/ethics as misaligned.
Taken together, we have:
Hypothesis.
Negative driver experiences which highlight job-fit through employee-firm interest alignment increase employee retention. Negative experiences where interests are not aligned decrease retention.
@@ 4.1 Data
Our empirical analysis uses communication records from a human resource outsourcing company specializing in retention programs for trucking companies. Primarily designed for new drivers, motor carriers typically apply these program for all new hires from training and orientation to three or six months after, depending on the program. [^3] Program details are determined by carriers. Our hazard models use fixed effects and flexible hazard function to control for heterogeneity across carriers, call center agents, and programs. See section 4.2 for further details. The core of the programs consists of weekly (or longer interval) phone calls to truck drivers to identify and record their “issues”– “shocks” in the terminology of the unfolding model of turnover– and then pass this information to their customers (i.e., trucking companies) to help them to be proactive in driver retention. Specifically, during the program, an agent of the human resource company records a brief summary of the driver conversation after each call, tagging it with proprietary classification codes for the issues identified during the conversation (hereafter issue codes). Some examples of the communication notes are below:
@@
The classification system includes more than two hundred issue codes that have nested structure allowing for coarse, intermediate and granular characterization of issues. [^4] Due to an non-disclosure agreement with the human resource company, the full list of issue codes and their definitions cannot be shared. For example, a non-competitive pay rate is classified as “Pay: Rate: Non-competitive,” while an unclear mileage rate is classified as “Pay: Rate: Confusing/Unclear.” In many cases, coarse issue codes share common cross-cutting intermediate categories, as when Covid issues show up as sub-categories in Pay, Training, Personal, etc.,. These cross-cutting categories allow masking of proprietary trade secrets by re-organizing the data without losing important issue-code variation. We re-classify primitive issue codes into the fourteen major categories of section 3.2: Home-time, Pay, Equipment, Safety/Ethics, Training, Team, Operations, Disrespect, Company, Covid, Waiting, Personal, Health, Others. In our main analysis, we further aggregate the categories into two grand classifications: “Aligned,” where employee-employer interests are aligned, including Equipment, Operations and Waiting, and “Misaligned,” where their interests are opposed, which includes all other categories.
@@
We merge these communication records with driver employment data to create survival-time data suitable for duration analysis. The employment data includes hire and termination dates, from which we calculate employment duration (in weeks) and termination status at the time of each communication record. For drivers whose employment was not terminated during the program, no termination date is observed and their employment spells are treated as censored (i.e., right-censoring). To create a more homogeneous sample, we exclude non-voluntary termination cases (i.e., disqualified quit) and employment spells longer than 100 weeks. [^5] Our results in the next sections are robust to this sample selection. The data is then merged with employer data such as carrier-level scores related to unsafe driving, hours-of-service compliance, vehicle maintenance, and misuse of controlled substances/alcohol from Safety Measurement System (SMS) of the Federal Motor Carrier Safety Administration.
@@
The duration and termination status variables, along with the communication records, are the key variables in our duration analysis. Our data is multi-record-per-subject survival data, where individual drivers are observed multiple times with different communication records over their employment spells. This provides our analysis additional strength because it allows for time-varying regressors that can account for the entire history of covariates. This feature is particularly useful for analyzing the effects of lagged variables.
@@
Our final sample includes 466,236 communication records (= total observations) of 45,873 employment spells from 21 carriers over 2019 - 2022. Most spells are from large carriers, with the largest carrier accounting for 60% and other seven carriers accounting for 32.6% of the sample. [^6] Employment concentration at larger firms is typical in this industry. While the majority of firms are small, the power-law distribution of firms size dictates that the majority of workers work for larger firms (Balthrop, 2021). Later years includes more observations: 16,882 (2019), 71,839 (2020), 158,221 (2021), 219,294 (2022). [^7] These are the numbers of communication records from the spells initiated in a given year. The sample includes four different retention programs, four call-intervals, and fifteen agents, where two programs account for 97.8% of the sample and most observations are based on weekly or bi-weekly phone calls (99.8%). Our duration analysis controls for heterogeneity across carriers, years, months, programs, call-intervals and agents using fixed effects and flexible hazard functions. [^8] In practice, the retention programs and call-intervals are determined by carriers, and in the data, program type and call-interval do not vary much within carriers (i.e., each carrier applies the same program and call-interval to all their drivers). This indicates that controlling for carrier-level heterogeneity is particularly important and an effective way to control for overall heterogeneity in our data.
@@
Next, we provide Kaplan-Meier (KM) survival curves and smoothed hazard estimates to summarize the important features in the data. The survival and hazard curves in Figures 1 and 2 plot survival rates, i.e., the proportion of employment spells that did not terminate after a given number of weeks, and hazard rates, i.e., the proportion of employment spells that terminated at a given week among those not yet terminated by that time, for each employment week. The shaded areas represent 95% confidence intervals.
@@
Figure 1 includes survival curves by carrier and year, highlighting important heterogeneity in the data. The carriers are ranked in terms of the number of employment spells and their rankings are denoted by “Top-.” The survival curves demonstrate the high turnover rate in trucking industry. Top1 and Top3-8 carriers’ survival rates are below 0.5 after 24 weeks, indicating that more than 50% of the drivers in these carriers quit their jobs within six months. Significant heterogeneity across carriers is observed. The survival rate of Top2 carrier is significantly higher than the other carriers, with the gap widening as the employment duration increases. The heterogeneity across years is also interesting. The survival curve of year 2019 flattens around fifteenth week, implying that driver turnover significantly decreases after the period. However, after 2020, turnover worsens, even for long-duration employees, providing evidence consistent with the trend of industry-wide employment re-shuffling in the wake of the Covid pandemic (Phares et al., 2023).
[$=$ Figure 2 here $=$]
Figure 2 presents survival and hazard curves for four different driver groups: i) the group of drivers experiencing no issues throughout their spells (26,777 spells), ii) the group experiencing Aligned issues only (6,402 spells), iii) the group experiencing Misaligned issues only (6,715 spells), and iv) the group experiencing both Aligned and Misaligned issues (5,979 spells). [^9] The hazard curves are smoothed by a kernel method in Klein & Moeschberger (2003). Note that the levels and curvatures of the plots do not directly represent the effects of the different types of issues since they do not account for the timing of the issues and heterogeneity across drivers. We account for these features in our duration models in the next section when calculating the marginal effects of the issues.
There is a significant difference in survival and hazard rates between groups experiencing Aligned issues and those not: groups experiencing Aligned issues exhibit higher survival rates and lower hazard rates over the whole employment period. Looking at the hazard curves, the discrepancy is pronounced in the early period, leading to a significant gap in survival rates in later periods. In contrast, the differences in survival and hazard rates between “Misaligned only” and “No issues” is quite small; similarly, “Aligned & Misaligned” closely tracks “Aligned only.”
@@ 4.2 Empirical Model: Cox Proportional Hazard Model
We use fixed effects and flexible hazard functions to control for heterogeneity across carriers, years, months, programs, call-intervals and agents. In all specifications, we include fixed effects for year, month, retention program, call-interval and agent. Also, we include a variable that quantifies the non-responsiveness of each driver to the human resource company’s calls, “Nonresponsive”, which is the proportion of calls tagged with “voice mail,” “no answer,” etc. Responsiveness is an important revealed characteristic that can be correlated with driver heterogeneity. The Nonresponsive variable is quite uniformly distributed across drivers, with quartiles at 0.20, 0.42, and 0.64, implying that the variable is not concentrated within a subset of individuals. To draw robust conclusions, standard errors are always clustered at the spell level.
Since much of the retention program’s details (length, call-interval, etc) are determined by carriers, and carrier characteristics are arguably the most important factors in driver’s quit decision, we employ three different approaches to control for heterogeneity across carriers: fixed effects, carrier-level variables from the SMS, and carrier-specific baseline hazard functions. In the result tables in the next sections, Model 1 and 2 includes carrier fixed effects and carrier-level variables, respectively, and Model 3 allows for carrier-specific baseline hazard functions. [^14] The partial likelihood estimation of the Cox PH model calculates the likelihood of each (observed) termination event by the hazard rate of the spell terminated at the event divided by the sum of the hazard rates of the terminated and non-terminated (i.e., censored) spells existing at the event. The division in the calculation cancels out the baseline hazard function due to the specification in (1), which allows us to estimate $\beta$ without specifying the baseline hazard function. When group-specific baseline hazard functions are allowed, the “partial” likelihoods are calculated by group.
@@ 5.1 Effects of Issue Types on Driver Turnover
To examine our hypothesis, Table 2 reports the coefficient estimates for Aligned and Misaligned issues in the Cox PH model. [^15] Remember that the estimate for each issue type is the average of the estimates for its original issue dummy and four lags, so it includes the dynamic effects. As defined in Section 4.1, Aligned, where employee-employer interests are aligned, includes Equipment, Operations and Waiting, and Misaligned, where their interests are opposed, includes all other categories. Coefficients estimates are stable across specifications.
@@
A stark contrast between the two types of issues is observed: all other things being equal, Aligned issues on average decrease driver turnover, whereas Misaligned issues increase it. “h.r.” are hazard ratios, which helps the reader quickly appreciate the magnitude of the impact. For example, in Model 1, the hazard ratio of Aligned is $-0.23$, indicates that experiencing a shock that highlights aligned interests decreases the hazard (i.e., instantaneous probability of leaving the job) by 23% over the baseline hazard. On the contrary, the hazard ratio for Misaligned issues is $0.40$, indicates having an issue like Home-time, Pay, etc., increases the hazard by nearly 40%. The somewhat counter-intuitive result here is that Aligned issues lower driver turnover, even though they are still negative events that drivers experience. For example, a driver that experiences problems with having to wait at a loading dock is more likely to remain at the motor carrier than a driver that experiences no wait-time shock.
We argue that this is because the Aligned issues provide opportunities for both drivers and employers to communicate with positive messages and confirm their mutual interests, thus increasing the employee’s embeddedness. When the driver experiences operational difficulty, equipment failure, or extended wait times, these shocks also harm the firm by directly reducing worker and equipment utilization. A worker detained at a loading dock, or broken down on the roadside, is not transporting freight. In these cases the motor carrier as an organization has incentive to sympathize with workers, which may result in strengthened relational linkages between frontline workers (drivers) and their support teams. Beyond words, the firm is also likely to take action to remedy the problem. Whether through words (which we observe) or actions (unobserved) these Aligned shocks highlight the firm and workers’ common purpose, while potentially revealing the firm’s support infrastructure available to the worker.
We further disaggregate these results in Table 3, which reports the coefficient estimates for fourteen major issue types in the Cox PH model. [^16] More detailed results including the estimates of individual lag terms are available upon request.
@@
As with our main results, coefficient estimates in Table 3 are not sensitive to specification, whether carrier fixed effects (Model 1), carrier-level variables (Model 2) or carrier-specific baseline hazards (Model 3) are included. Categories most strongly associated with increased turnover include Covid and Personal issues. While Covid is a strong indicator of potential exit, it is a comparatively rare issue (only 224 instances), even though our sample period covers the onset of the pandemic. Personal issues are more common, and have stronger effects on turnover. This category encompasses issues related to family, which supports previous research establishing loneliness and family disruption as one of the most important reasons for truck driver turnover (Williams et al., 2017). Even the smaller (but positive) coefficients have managerially meaningful effects on hazard: ethics complaints increase the baseline hazard rate by 23% (Model 1). In contrast, the Aligned issues, including Equipment, Operations, and Waiting, all result in strongly reduced hazard rates. The coefficient on Nonresponsive variable (i.e., proportion of calls not answered) is reported in Table 2, which is, as expected, positive and significant, with estimates ranging from 0.63 to 0.67 across models, implying drivers less responsive to the retention programs exhibit higher turnover rates.
5.2 Robustness Check
@@
A potential concern, however, is that there may be unobserved driver types with different hazard rates and these types are correlated with the kinds of issues they report. For example, there may be “attached” and “detached” drivers with low and high turnover rates, and the former report aligned issues more frequently than the latter, simply because they are more attached to their company. Then, the signs of the coefficients on Aligned and Misaligned issues are simply a reflection of the hazard rate difference between the two types, not the effects of the two different types of issues. Said differently, it is possible to imagine a driver experiencing an equipment issue but not reporting it precisely because the driver expects the issue will go unresolved. Such a disillusioned driver can be expected to have a higher turnover rate; importantly, this type of detached driver will fall into the base-case “no issue” category, biasing base-case turnover rates upward. The Nonresponsive variable in our models accounts for the issue in part, but this section explores various approaches to further examine the robustness of our results. We pursue three strategies: i) controlling for additional covariates, ii) focusing on a sub-sample of responsive drivers, and iii) examining for differences between employee drivers and owner-operators.
First, we include “issue frequency” variables into our models and report the results in Table 4. Issue frequencies are calculated for Aligned and Misaligned issues, each frequency being the number of each driver’s Aligned or Misaligned issues reported throughout their sample period divided by the period length (hereafter Aligned_f and Misaligned_f). To account for potential non-linearity, we also include their squared terms. The frequency variables essentially allow us to compare outcomes among those who exhibit similar frequencies of Aligned or Misaligned issues throughout their employment spells.
After inclusion of the variables, the signs and significance of Aligned and Misaligned variables remain the same. However, the two coefficients exhibit somewhat different patterns. The coefficient on Misaligned issues decreases significantly (about 35%), while the coefficient on Aligned issues largely remains the same. The result can be interpreted as follows. Recall that our model includes dummies for Aligned and Misaligned issues and their lags up to four periods (as stated in Footnote 11, we used the LASSO to determine the optimal number of lags). The coefficient estimate for Aligned or Misaligned issues reported in this paper is the average of the estimates for its original dummy and four lags. Therefore, the coefficient estimate roughly represents the likelihood of employment termination “within” five weeks after Aligned or Misaligned issues. In this setup, the frequency variables are likely to capture turnovers beyond five weeks. Therefore, the decrease in the coefficient on Misaligned indicates that there were many cases where drivers did not leave the company within five weeks after Misaligned issues but eventually left the company (without any issues reported). The coefficient on Aligned issues remains the same, implying there is no such pattern for Aligned issues. [^17] As discussed below, the coefficients on Aligned_f and (Aligned_f)2 in Table 4 are opposite (negative and positive), implying a nonlinear effect of Aligned_f, which makes the overall effect to be close to zero.
These turnovers may be seen as a long-term effect of Misaligned issues but can also be seen as a result of “persistent” driver heterogeneity that might have affected both issue reporting behavior and turnover. Indeed, this approach is similar to an econometric treatment popular in practice for accounting for individual heterogeneity: the correlated random effects (CRE) approach (e.g., Wooldridge, 2019). The CRE approach specifies individual heterogeneity as a function of the time-averages of the covariates in the model and includes these additional variables into the model to control for the heterogeneity. The inclusion of frequency variables can therefore be viewed as an application of the CRE approach, with the results in Table 4 demonstrating the robustness of our conclusion to individual heterogeneity.
The coefficients on the frequency variables and their squared terms deserve close examination. The coefficient on Aligned_f is negative but the coefficient on its squared term (i.e., (Aligned_f)2) is positive, implying a non-linear relationship between the number of Aligned issues and turnover rate. For those with relatively low Aligned issue frequencies, experiencing an additional Aligned issue decreases turnover, but when the frequencies are high, it increases turnover. Drivers reporting issues very frequently may be seen as “detached” drivers, regardless of issue type, and the squared term captures their inherently high hazard rates. The coefficients on Misaligned_f and its squared term are all positive, indicating higher frequency is always associated with higher turnover for Misaligned issues.
@@
As a second strategy, we focus on the sub-population of observed communicators. That is, rather than making comparison to a reference group that may under-report issues, we make comparisons between groups that report multiple issues. First, we dis-aggregate the Aligned category into “Broken-equipment” and “Other aligned,” where “Broken-equipment” includes issues related to “broken/missing components” only, which is arguably less subject to the under-reporting problem. Second, we re-classify the issues into new categories that account for joint-distribution of Aligned and Misaligned issues: Broken-equipment only (B-only), Other aligned only (OA-only), Misaligned only (MA-only) and their combinations. [^18] Note that a communication record can be tagged with multiple codes in different categories. Communication records counted multiple times for the fourteen categories in Table 1 can be counted fewer times for the new categories in Table 5, resulting in different total numbers in the tables.
@@
We note that, if other covariates are ignored, the coefficient on Aligned in Table 2 is calculated by a combination of two hazard rate differences: a hazard rate difference between the cases experiencing Aligned issues only and the cases experiencing no issues, and another hazard rate difference between the cases experiencing both Aligned and Misaligned issues and the cases experiencing Misaligned issues only. Roughly speaking, the first difference compares drivers reporting Aligned issues only and drivers who were silent, whereas the second difference compares drivers who report at least one of the two types of issues. Because the second comparison is among “communicators,” it is arguably less subject to the under-reporting issue. The new categories allow us to identify these dis-aggregated effects.
@@
Table 5 includes the distribution of the issues under the new categories, and Table 6 reports the coefficient estimates for the new categories. The coefficient on “Broken & Misaligned” subtracted by the coefficient on “Misaligned” calculates the effect of Aligned issues that is less subject to the under-reporting issue as argued above, which is -0.127 in Model 1 (i.e., specification including carrier fixed effects), with standard error 0.078, t-statistic -1.63 and p-value 0.10.191919The standard error is calculated using the delta method. The results from Model 2 and 3 specifications are qualitatively the same. The p-value is 10%, but the result overall indicates that relative to the population experiencing only Misaligned shocks, a driver experiencing both an Aligned and Misaligned shock has a lower hazard rate. The striking result that two negative shocks are better than one shock provides strong support for our hypothesis, at least in this sub-population.
@@
Third, we examine how the results change when we account for equipment ownership status. Carriers may own their tractors or lease them by contracting with owner-operators. Owner-operators are responsible for the maintenance of their own tractors; therefore, we can expect smaller effects of equipment-related shocks in this sub-population. Unfortunately, our data doesn’t include information on individual-level equipment ownership, so we use carrier-level lease rates for this exercise. We calculate lease rates for each carrier by the proportion of leased tractors, [^20] The average lease rate for 21 carriers over the years 2019-2021 is 21.02%. using data from the FMCSA Census. We select two sub-samples containing spells from the carriers with the average lease rate higher than 40% over the years 2019-2021 (“Lease$>$ 40”) and from those with the lease rate lower than 5% (“Lease$<$ 5”). Each sub-sample accounts for about 10% of the whole sample. [^21] We use these cutoffs to select two sub-samples that are comparable in terms of size and reasonably large to get reliable results.
Table 7 report the results from the two sub-samples using the Model 3 specification in Table 6. The results from other specifications are qualitatively the same. We also report the distribution of the issue codes in each sub-sample in Table 8. The sub-sample “Lease$>$ 40” contains fewer Broken-equipment only shocks and more Misaligned shocks than “Lease$<$ 5”, but the overall distribution of the issues is similar across the two sub-populations, enhancing the comparability of the results in Table 7 across the two samples.
[$=$ Table 7 here $=$]
[$=$ Table 8 here $=$]
For Lease$>$ 40 (i.e., the sample including many owner-operators), most of the coefficients related to equipment shocks are now insignificant. It appears that only Misaligned and Other aligned & Misaligned are significant predictors for driver turnover in this sample. This aligns with our prediction that if the equipment-related coefficients indeed capture the effects of equipment shocks, not only reflecting driver heterogeneity or inherent hazard rate differences between driver types, the effects will be smaller when drivers own and maintain their equipment. The results from Lease$>$ 40 are similar to those in Table 6.